ue or false. MPC+ MPS 1 o True O False Correct. Marginal propensity to consume plus marginal propensity to save equals 1. of aggregate expenditure is 5 MPcio 8 when the MPC is.8 and there is an increase in investment spending of $100,000. x Incorrect. First determine the expenditure multiplier, then multiply that by $100,000 to obtain the correct answer True or false. If people save more of their income, the expenditure multiplier will not decrease and aggregate expenditures will decrease True False MacBook Pro
Answer #1

Q1.

Answer: $500,000

Expenditure multiplier = 1 / (1 – MPC)

                                    = 1 / (1 – 0.8)

                                    = 1 / 0.2

                                    = 5

Required value of AE = Expenditure multiplier × Investment

                                    = 5 × $100,000

                                    = $500,000

Q2.

Answer: False

Take an example for understanding the fact.

Suppose MPC (as above) is 0.8 and MPS is 0.2; therefore, AE becomes $500,000.

Now if MPS becomes 0.4, multiplier would be (1 / 0.4 =) 2.5; therefore, AE would be (2.5 × $100,000 =) $250,000.

Conclusion: If there is more saving, multiplier and AE will decrease certainly.